OpenStudy (anonymous):

Simplify 3/a + 6/a^2 - 9/a^3

OpenStudy (anonymous):

\[3/a + 6/a^2-9/a^3\]

OpenStudy (anonymous):

Put them together with a^3 as a common denominator \[\frac{3 a^2+6 a-9}{a^3} =\frac{3 \left(a^2+2 a-3\right)}{a^3} \] Can you still factor the numerator?

OpenStudy (anonymous):

Notice that the numerator can be factored as 3 (-1 + a) (3 + a) Hence the final nice form is \[ \frac{3 (a-1) (a+3)}{a^3} \]